A wooden cube is found to float in water with 1/2cm of of its vertical side above the water.on keeping a weight of 50gm over its top, it is just submerged in the water find the specific gravity

Grade:12

2 Answers

Utsav Basu
6 years ago
Let the side of the cube be a cm
Thus the volume of above water surface = (a2/2)cm3
So, a volume of  (a2/2)cm3 of the cube is sufficient to displace 50gm i.e. 50cm3 of water to support the 50gm place upon it.
Thus, (a2/2)cm3 = 5cm3
or,      a = 10cm
 
Therefore, after placing the weight of 50gm,
Buoyant force = 1000g
Weight = (m + 50)g
 
But, Buoyant force = Weight
So, 1000g = (m + 50)g
or,  m = 950
 
Thus, density of cube = 950gm/1000cm3 = 0.95gm/cm3
Therefore, specific gravity = 0.95gm/cm3
HABEEB UR RAHMAN
4 years ago
 
 
Let the side of the cube be s cm
Thus the volume of above water surface = (s2/2)cm3
So, a volume of  (s2/2)cm3 of the cube is sufficient to displace 50gm i.e. 50cm3 of water to support the 50gm place upon it.
Thus, (s2/2)cm3 = 5cm3
or,      s= 10cm
 
Therefore, after placing the weight of 50gm,
Buoyant force = 1000g
Weight = (m + 50)g
 
But, Buoyant force = Weight
So, 1000g = (m + 50)g
or,  m = 950
 
Thus, density of cube = 950gm/1000cm3 = 0.95gm/cm3
Therefore, specific gravity = 0.95gm/cm3